Recently, More Money 4U offered an annuity that pays 4.8% compounded monthly. If $1,000 is deposited into this annuity every month, how much is in the account after 6 years? How much of this is interest?

Type the amount in the account: $____

Type the amount of interest earned: $____

(Please square, label, or circle the answer)
Answer #1

Solution:

Monthly interest rate = 4.8%/12 = 0.40%

Monthly deposit = $1,000

Nos of monthly periods = 6*12 = 72 periods

Amount in the account after 6 years = $1,000 * Cumulative FV factor at 0.40% for 72 periods

= $1,000 * 83.24783 = $83,248

Amount of interest earned = $83,248 - $72,000 = $11,248